Find the solution by substitution:
3x-2y=-5
x=y-2

Answer by checkley77(12844)

3x-2y=-5
x=y-2
3(Y-2)-2Y=-5
3Y-6-2Y=-5
Y=6-5
Y=1 ANSWER.
X=1-2
X=-1 ANSWER.
PROOF:
3*-1-2*1=-5
-3-2=-5
-5=-5